OpenStudy (anonymous):

Simplify the expression below. (x^6)(x^2) x^4 X^6 X^8 X^12

OpenStudy (acxbox22):

\[(x^a)(x^b)=x ^{a+b}\]

OpenStudy (anonymous):

so would 4^5 x 4^3= 4^8?

OpenStudy (anonymous):

@acxbox22 would 4^5 x 4^3= 4^8?

OpenStudy (acxbox22):

yes